# Fast Waveform Generation for Gravitational Waves using Evolutionary Algorithms

### v1

@@ Abstract
Gravitational-wave analyses depend heavily on waveforms that model the evolution of compact binary coalescences as seen by observing detectors. In many cases these waveforms are given by waveform approximants, models that approximate the power of the waveform at a set of frequencies. Because of their omnipresence, improving the speed at which approximants can generate waveforms is crucial to accelerating the overall analysis of gravitational-wave detections. An optimisation algorithm is proposed that can select at which frequencies in the spectrum an approximant should compute the power of a waveform, and at which frequencies the power can be safely interpolated at a minor loss in accuracy. The algorithm used is an evolutionary algorithm modeled after the principle of natural selection, iterating frequency arrays that perform better at every iteration. As an application, the candidates proposed by the algorithm are used to reconstruct signal-to-noise ratios. It is shown that the IMRPhenomXPHM approximant can be sped up by at least $30\%$ at a loss of at most $2.87\%$ on the drawn samples, measured by the accuracy of the reconstruction of signal-to-noise ratios. The behaviour of the algorithm as well as lower bounds on both speedup and error are explored, leading to a proposed proof of concept candidate that obtains a speedup of $46\%$ with a maximum error of $0.5\%$ on a sample of the parameter space used.
@@ II Approximants
Since this speed-up is obtained through interpolation and not algorithmically, $h$ and $h^{*}$ can not be expected to be equal, and a loss in accuracy will be incurred. Two metrics will be used to measure this loss. The first is the residual sum of squares for the real and imaginary parts of $h$ and $h^{*}$. This metric is used by the evolutionary algorithm internally. The second metric is based on a widely used quantity, the signal-to-noise ratio (SNR).
@@ IV.2 Hyperparameters
Due to a symmetry argument applied to the phenotype construction process, the algorithm can be ran both forwards and backwards. In the forwards case, the phenotype grows in size towards the full frequency spectrum as frequencies are added in each iteration of the algorithm. In the backwards case the frequency spectrum is reduced in size towards the background by removing frequencies in each iteration. As a rule of thumb the forwards case is better suited to finding smaller frequency arrays, whereas the backwards case is better suited to finding frequency arrays that result in a low error. Note that the parameter that dictates the direction stands in direct relation to the penalty $w$ included in Eq. 5. In both cases the algorithm needs to be incentivised to move through the solution space by the penalty. In for instance the backwards direction, if $w$ were to be set to 00, there would be no reason to remove any frequency from the phenotype, as doing so would only lead to a loss in accuracy. Setting $w>0$, the removal of this same frequency might result in an increased fitness, as long as the loss in accuracy is less than the penalty $w$ for including the frequency.
@@
Recapitulating this subsection, besides the initial choices of background size during pre-processing and the probability measure for the parent selection function, the algorithm is defined by the hyperparameters $(\delta,w,g,d,s,r)$ given in Table 1. The values in this table represent the hyperparameters used in this proof of concept, with True for the direction $d$ meaning the algorithm was ran backwards. This choice was made to prioritise accuracy, which also inspired the choice of $\delta$. The runtime $r$ was set to well above the observed limit of stagnation in trial runs, with the population size $g$ and frequency sample size $s$ empirically determined following the previous discussion. The penalty weight $w$ is set to $8$, but any value to counterbalance the increasing modulus in the objective function as the evolutionary algorithm progresses and frequencies are removed would have sufficed. These choices of hyperparameters consistently drive the evolutionary algorithm to converge to similar and well-performing frequency array phenotypes on repeated runs, suggesting the optimisation does not get stuck in easily avoidable local minima. The runtime for the algorithm (including pre-processing) on an Intel Gold 6154 CPU is in the order of magnitudes of minutes.
@@ IV.3 Performance measures
Finding the neighbourhood in the parameter space $\mathcal{P}$ where the frequency arrays are effective is important, as the reference event input to the algorithm during pre-processing adds a level of specificity. Nearby the reference events the frequency arrays are likely to allow a more accurate reconstruction of the SNR. In order to analyse the generalisibility, $10,000$ points are sampled from the default precessing binary black hole prior in Bilby version 2.2.2 (which can be seen in Table 3). At every sampled set of parameters, the exact plus-polarisation of the waveform is generated and injected into simulated noise generated using the LIGO Livingston PSD included in Bilby version 2.2.2. The SNR of both the exact plus-polarisation, hereafter also called the exact SNR, and the approximate plus-polarisation, hereafter called the approximate SNR or $\textup{SNR}^{*}$, is computed against this strain. The percentual error for the frequency array $F$ is then calculated as:
@@ V Results
Table 3: The ranges for each of the variables in the parameter space and the recovered ranges per frequency array. If no unit is included, the quantity is dimensionless. Note that the geocent time is not included, as it is inconsequential to this test. All parameter ranges save for the luminosity distance and chirp mass were recovered, with both of these not being fully covered by the drawn sample of parameters.
The results of the generalisibility tests to determine the regions of the parameter space $\mathcal{P}$ where the frequency arrays are effective are shown in Table 3. For each of the $15$ parameters (minus the geocent time) the sampled ranges are shown, as well as the ranges that were recovered for the frequency arrays. Recovered here means that the parameter samples from these ranges were accepted by their error being below the given threshold, with $E_{\%}[F]<0.01$. For most parameters, the full ranges were recovered, save for the luminosity distance and chirp mass. In the case of the luminosity distance, further inspection reveals that the lower ranges were simply not reached in the parameter sample drawn. The same holds for the chirp masses due to the way Bilby samples the chirp mass through uniform sampling in the component masses.
@@
Investigating the cause for the rejection of parameter samples, it becomes apparent that nearly all rejected samples have a low exact SNR value. This is visualised in Fig. 9, showing the exact SNR distributions for the rejected parameter samples versus the exact SNR distributions for the full samples. It can be seen from this figure that for the rejected parameter samples, the SNR is on average an order of magnitude smaller than that of the full parameter sample, and moreover that the distributions are left-tail heavy. In order to verify the strength of the SNR as a predictor for rejection, logistic regression Bishop (2006) of the binary acceptance class onto the waveform parameters was employed. The sampled parameters for the three candidates were stacked, with the luminosity distance replaced by the SNR. The removal of the luminosity distance is a requirement so that the multicollinearity assumption of logistic regression is not violated, as luminosity distance correlates strongly with SNR. Despite the SNR scaling with several other parameters, the removal of the luminosity distance should sufficiently reduce the correlation. A binary labelling was created for this dataset by assigning a dependent variable $1$ to the accepted parameter samples and 00 to the rejected samples. Fitting a logistic regression model, the SNR was found to be the predictor with the highest coefficient by far, with a value of $\beta_{\textup{SNR}}=16.5$. This coefficient was followed by the chirp mass with a value of $\beta_{\textup{Chirp}}=3.74$, meaning the SNR is a significantly stronger predictor for the rejection of a parameter sample according to the regression model. As with accuracy, the generalisibility too is bounded from below by the results presented.
@@
Based on the analysis in this section, the frequency array obtained by using GW150914 as the reference event is proposed as a proof of concept, as it obtains a speed-up in computational time of $46\%$ at a maximum error of $0.5\%$ in SNR. A heatmap for the generalisibility of this frequency array is shown in Fig. 10. This figure was created by drawing $10$ partial parameter samples, meaning every parameter except for the mass ratio and chirp mass were sampled, and then varying the mass ratio and chirp mass before computing the errors $E_{\%}$ for every such combination. A point in the plane then corresponds to the average error of the $10$ parameter samples for that combination of the chirp mass and mass ratio, and the full figure shows $100,000$ such parameter samples in total. It can be seen from the figure that especially in the band of the chirp mass where the event belongs to, moving away from the mass ratio of the event will increase the error.
@@
for $h$ the waveform projected onto the detector frame and $D$ the strain. Setting $D:=N+h^{\prime}$ for $N$ noise, the maximum of the log likelihood should be attained at the injected waveform $h^{\prime}$. Analogous to the SNR reconstruction, the exact log likelihood maximum $\mathcal{L}(h,N+h)$, or $\mathcal{L}_{m}$ for shorthand, and approximate log likelihood maximum $\mathcal{L}(h^{*},N+h)$, or $\mathcal{L}^{*}_{m}$, were computed for a parameter sample of size $10,000$. Here, however, both polarisations are included. Note that these errors do not need to be converted to percentages as was done for the errors in SNR, since the variance in the values will be significantly smaller. The absolute error distribution is shown in Fig. 11 and shows promising results. The reader should be aware however that the evolutionary algorithm was not finetuned to this computation, suggesting that a higher accuracy can likely be obtained. Specifically, the objective function should be adapted to the log likelihood function, and the hyperparameters should be rebalanced.
Further study in general might reveal marginally better performing frequency arrays, or frequency arrays better suited to specific sections of the parameter space. It is however expected that the improvement from frequency subsampling and linear interpolation represented by the values in this section is near-optimal. Implementing the naive frequency subsampling scheme of interpolating every second frequency (setting $\rho=0.5$) for GW150914, the speedup $S$ evaluates to $1.99$, but the minimum error of $E_{\%}$ blows up by a factor with order of magnitude $10^{4}$, in the process causing a rejection for 5972 parameter samples.
@@ Acknowledgements
With thanks to Harsh Narola, Stefano Schmidt, Melissa Lopez, Justin Janquart, Bhooshan Gadre, Marc van der Sluys, Chris van den Broeck and Justin Perez. Q.M is supported by the research program of the Netherlands Organisation for Scientific Research (NWO). S.C is supported by the National Science Foundation under Grant No. PHY-2309332. The authors are grateful for computational resources provided by the LIGO Laboratory and supported by the National Science Foundation Grants No. PHY-0757058 and No. PHY-0823459. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation.

### v2

@@ Abstract
Gravitational-wave analyses depend heavily on waveforms that model the evolution of compact binary coalescences as seen by observing detectors. In many cases these waveforms are given by waveform approximants, models that approximate the amplitude and phase of the waveform at a set of frequencies. Because of their omnipresence, improving the speed at which approximants can generate waveforms is crucial to accelerating the overall analysis of gravitational-wave detections. An optimisation algorithm is proposed that can select at which frequencies in the spectrum an approximant should compute the power of a waveform, and at which frequencies the power can be safely interpolated at a minor loss in accuracy. The algorithm used is an evolutionary algorithm modeled after the principle of natural selection, iterating frequency arrays that perform better at every iteration. As an application, the candidates proposed by the algorithm are used to reconstruct signal-to-noise ratios. It is shown that the IMRPhenomXPHM approximant can be sped up by at least $30\%$ at a loss of at most $2.87\%$ on the drawn samples, measured by the accuracy of the reconstruction of signal-to-noise ratios. The behaviour of the algorithm as well as lower bounds on both speedup and error are explored, leading to a proposed proof of concept candidate that obtains a speedup of $46\%$ with a maximum error of $0.5\%$ on a sample of the parameter space used.
@@ II Approximants
Since this speed-up is obtained through interpolation and not algorithmically, $h$ and $h^{*}$ can not be expected to be equal, and a loss in accuracy will be incurred. Two metrics will be used to measure this loss. The first is the residual sum of squares (RSS) for the real and imaginary parts of $h$ and $h^{*}$. This metric is used by the evolutionary algorithm internally. The second metric is based on a widely used quantity, the signal-to-noise ratio (SNR).
@@ IV.2 Hyperparameters
Due to a symmetry argument applied to the phenotype construction process, the algorithm can be ran both forwards and backwards. In the forwards case, the phenotype grows in size towards the full frequency spectrum as frequencies are added in each iteration of the algorithm. In the backwards case the frequency spectrum is reduced in size towards the background by removing frequencies in each iteration. As a rule of thumb the forwards case is better suited to finding smaller frequency arrays, whereas the backwards case is better suited to finding frequency arrays that result in a low error. This is a result of the stopping conditions typically being met before the frequency array reaches half the size of the frequency spectrum. Note that the parameter that dictates the direction stands in direct relation to the penalty $w$ included in Eq. 5. In both cases the algorithm needs to be incentivised to move through the solution space by the penalty. In for instance the backwards direction, if $w$ were to be set to 00, there would be no reason to remove any frequency from the phenotype, as doing so would only lead to a loss in accuracy. Setting $w>0$, the removal of this same frequency might result in an increased fitness, as long as the loss in accuracy is less than the penalty $w$ for including the frequency.
@@
Recapitulating this subsection, besides the initial choices of background size during pre-processing and the probability measure for the parent selection function, the algorithm is defined by the hyperparameters $(\delta,w,g,d,s,r)$ given in Table 1. The values in this table represent the hyperparameters used in this proof of concept, with True for the direction $d$ meaning the algorithm was ran backwards. This choice was made to prioritise accuracy, which also inspired the choice of $\delta$. The runtime $r$ was set to well above the observed limit of stagnation in trial runs, with the population size $g$ and frequency sample size $s$ empirically determined following the previous discussion. The penalty weight $w$ is set to $8$, but any value to counterbalance the increasing modulus in the objective function as the evolutionary algorithm progresses and frequencies are removed would have sufficed. These choices of hyperparameters consistently drive the evolutionary algorithm to converge to similar and well-performing frequency array phenotypes on repeated runs, suggesting the optimisation does not get stuck in easily avoidable local minima. The runtime for the algorithm (including pre-processing) on an Intel Gold 6154 CPU is in the order of magnitudes of minutes, bound by the longest recorded run.
@@ IV.3 Performance measures
Finding the neighbourhood in the parameter space $\mathcal{P}$ where the frequency arrays are effective is important, as the reference event input to the algorithm during pre-processing adds a level of specificity. Nearby the reference events the frequency arrays are likely to allow a more accurate reconstruction of the SNR. In order to analyse the generalisibility, $10,000$ points are sampled from the default precessing binary black hole prior in Bilby version 2.2.2. At every sampled set of parameters, the exact plus-polarisation of the waveform is generated and injected into simulated noise generated using the LIGO Livingston PSD included in Bilby version 2.2.2. The SNR of both the exact plus-polarisation, hereafter also called the exact SNR, and the approximate plus-polarisation, hereafter called the approximate SNR or $\textup{SNR}^{*}$, is computed against this strain. The percentual error for the frequency array $F$ is then calculated as:
@@ V Results
In the current implementation, the error is more pronounced in the phase than it is in amplitude. Depending on the application, the objective function can be modified to either balance this error, or favor one of the two.
@@
As a note on data compression and memory usage, before interpolating to the full frequency spectrum, one can opt to write the waveform to disk, indexed by the frequency array. By choosing to interpolate to the full spectrum later, waveforms can be stored in a compressed format, and the compression factor will equal $\rho$.
@@
Table 3: The ranges for each of the variables in the parameter space and the recovered ranges per frequency array, in rounded numerical format. If no unit is included, the quantity is dimensionless. Note that the geocent time is not included, as it is inconsequential to this test. All parameter ranges save for the luminosity distance and chirp mass were recovered, with both of these not being fully covered by the drawn sample of parameters.
The results of the generalisibility tests to determine the regions of the parameter space $\mathcal{P}$ where the frequency arrays are effective are shown in Table 3. For each of the $15$ parameters (minus the geocent time) the sampled ranges are shown, as well as the ranges that were recovered for the frequency arrays. Recovered here means that the parameter samples from these ranges were accepted by their error being below the given threshold, with $E_{\%}[F]<0.01$. For most parameters, the full ranges were recovered, save for the luminosity distance and chirp mass. In the case of the luminosity distance, further inspection reveals that the lower ranges were simply not reached in the parameter sample drawn. The same holds for the chirp masses due to the way Bilby samples the chirp mass through uniform sampling in the component masses. If this is manually changed the results might vary slightly.
@@
Investigating the cause for the rejection of parameter samples, it becomes apparent that nearly all rejected samples have a low exact SNR value for $h$. This is visualised in Fig. 9, showing the exact SNR distributions for the rejected parameter samples versus the exact SNR distributions for the full samples. It can be seen from this figure that for the rejected parameter samples, the SNR is on average an order of magnitude smaller than that of the full parameter sample, and moreover that the distributions are left-tail heavy. In order to verify the strength of the SNR as a predictor for rejection, logistic regression Bishop (2006) of the binary acceptance class onto the waveform parameters was employed. The sampled parameters for the three candidates were stacked, with the luminosity distance replaced by the SNR. The removal of the luminosity distance is a requirement so that the multicollinearity assumption of logistic regression is not violated, as luminosity distance correlates strongly with SNR. Despite the SNR scaling with several other parameters, the removal of the luminosity distance should sufficiently reduce the correlation. A binary labelling was created for this dataset by assigning a dependent variable $1$ to the accepted parameter samples and 00 to the rejected samples. Fitting a logistic regression model, the SNR was found to be the predictor with the highest coefficient by far, with a value of $\beta_{\textup{SNR}}=16.5$. This coefficient was followed by the chirp mass with a value of $\beta_{\textup{Chirp}}=3.74$, meaning the SNR is a significantly stronger predictor for the rejection of a parameter sample according to the regression model. As with accuracy, the generalisibility too is bounded from below by the results presented.
@@
Based on the analysis in this section, the frequency array obtained by using GW150914 as the reference event is proposed as a proof of concept, as it obtains a speed-up in computational time of $46\%$ at a maximum error of $0.5\%$ in SNR. A heatmap for the generalisibility of this frequency array is shown in Fig. 10. This figure was created by drawing $10$ partial parameter samples, meaning every parameter except for the mass ratio and chirp mass were sampled, and then varying the mass ratio and chirp mass before computing the percentual error in SNR $E_{\%}$ for every such combination. A point in the plane then corresponds to the average error of the $10$ parameter samples for that combination of the chirp mass and mass ratio, and the full figure shows $100,000$ such parameter samples in total. It can be seen from the figure that especially in the band of the chirp mass where the event belongs to, moving away from the mass ratio of the event will increase the error.
@@
for $h$ the waveform projected onto the detector frame and $D$ the strain. Setting $D:=N+h^{\prime}$ for $N$ noise, the maximum of the log likelihood should be attained at the injected waveform $h^{\prime}$. Analogous to the SNR reconstruction, the exact log likelihood maximum $\mathcal{L}(h,N+h)$, or $\mathcal{L}_{m}$ for shorthand, and approximate log likelihood maximum $\mathcal{L}(h^{*},N+h)$, or $\mathcal{L}^{*}_{m}$, were computed for a parameter sample of size $10,000$. Here, however, both polarisations are included. Note that these errors do not need to be converted to percentages as was done for the errors in SNR, since the variance in the values will be significantly smaller. The absolute error distribution is shown in Fig. 11 and shows promising results. The reader should be aware however that the evolutionary algorithm was not finetuned to this computation, suggesting that a higher accuracy can likely be obtained. Specifically, the objective function (that defines the optimisation problem) should be adapted to the log likelihood function, and the hyperparameters should be rebalanced.
Further study in general might reveal marginally better performing frequency arrays, or frequency arrays better suited to specific sections of the parameter space. It is however expected that the improvement from frequency subsampling and linear interpolation represented by the values in this section is near-optimal. Implementing the naive frequency subsampling scheme of interpolating every second frequency (setting the relative array size at $\rho=0.5$) for GW150914, the speedup $S$ evaluates to $1.99$, but the minimum error of $E_{\%}$ blows up by a factor with order of magnitude $10^{4}$, in the process causing a rejection for 5972 parameter samples.
@@ VI Conclusions
For the evolutionary algorithm itself, the investigation of its applications to data compression for gravitational-wave data analysis could prove useful. This can be coupled to the exploration of alternative interpolation schemes.
@@ Acknowledgements
With thanks to Harsh Narola, Stefano Schmidt, Melissa Lopez, Justin Janquart, Bhooshan Gadre, Marc van der Sluys, Chris van den Broeck, Justin Perez and the anonymous referee. Q.M is supported by the research program of the Netherlands Organisation for Scientific Research (NWO). S.C is supported by the National Science Foundation under Grant No. PHY-2309332. The authors are grateful for computational resources provided by the LIGO Laboratory and supported by the National Science Foundation Grants No. PHY-0757058 and No. PHY-0823459. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation.